OpenStudy (anonymous):

find the equation for the slope of a line for -2x+y=-4

OpenStudy (welshfella):

rearrange the equation to the form y = mx + c m is the required slope

OpenStudy (anonymous):

^ where y=y m=-2 x=x -4 = c

OpenStudy (welshfella):

do you know how to do thet?

OpenStudy (anonymous):

y=-2x+-4

OpenStudy (welshfella):

not quite you add 2x to both sides so what is the sign of 2x?

OpenStudy (anonymous):

Not exactly, using algebra you have to move the -2x to the other side other equation

OpenStudy (anonymous):

Srry I kind of confused you

OpenStudy (welshfella):

y - 2x + 2x = 2x - 4

OpenStudy (anonymous):

y=2x+-4 x=-2

OpenStudy (welshfella):

the 2x's on the left disappear to give y = ?

OpenStudy (anonymous):

y=2-4

OpenStudy (welshfella):

right y = 2x - 4 is correct and what is the slope of the line?

OpenStudy (anonymous):

-2x

OpenStudy (welshfella):

compare y = mx + c y = 2x - 4

OpenStudy (welshfella):

not - 2x - the slope is a number slope = m in the general equation

OpenStudy (welshfella):